# NoLiMa: Long-Context Evaluation Beyond Literal Matching

### v2

@@ Abstract
Recent large language models (LLMs) support long contexts ranging from 128K to 1M tokens. A popular method for evaluating these capabilities is the needle-in-a-haystack (NIAH) test, which involves retrieving a “needle” (relevant information) from a “haystack” (long irrelevant context). Extensions of this approach include increasing distractors, fact chaining, and in-context reasoning. However, in these benchmarks, models can exploit existing literal matches between the needle and haystack to simplify the task. To address this, we introduce NoLiMa, a benchmark extending NIAH with a carefully designed needle set, where questions and needles have minimal lexical overlap, requiring models to infer latent associations to locate the needle within the haystack. We evaluate 12 popular LLMs that claim to support contexts of at least 128K tokens. While they perform well in short contexts ($<$1K), performance degrades significantly as context length increases. At 32K, for instance, 10 models drop below 50% of their strong short-length baselines. Even GPT-4o, one of the top-performing exceptions, experiences a reduction from an almost-perfect baseline of 99.3% to 69.7%. Our analysis suggests these declines stem from the increased difficulty the attention mechanism faces in longer contexts when literal matches are absent, making it harder to retrieve relevant information. We publicly release the dataset and evaluation code at https://github.com/adobe-research/NoLiMa.
@@ 1 Introduction
We evaluate NoLiMa over 12 state-of-the-art language models, all claiming to support token lengths of at least 128K, including GPT-4o, Gemini 1.5 Pro, and Llama 3.3 70B (Hurst et al., 2024; Team et al., 2024a; Meta, 2024). Unlike NIAH-based evaluations, which contain literal matches and exhibit near-saturated performance, NoLiMa presents a more demanding challenge that highlights the limitations of these models. We evaluate 12 state-of-the-art language models, all of which claim to support token lengths of at least 128K, including GPT-4o, Gemini 1.5 Pro, and Llama 3.3 70B (Hurst et al., 2024; Team et al., 2024a; Meta, 2024). Their performance declines noticeably as context length increases, with considerable drops even at 2K–8K tokens. For instance, at 32K tokens, 10 out of 12 models achieve only half of their short-context performance.
@@
  • Impact of Latent Hops and Fact Direction: We demonstrate how the number of associative reasoning steps (latent hops) and the ordering of elements within a fact statements influence task performance.
  • Context Length vs. Needle Position: Our aligned-depth analysis shows that as latent reasoning complexity grows, performance depends more on context length than needle position. Without surface-level cues, longer contexts overwhelm the attention mechanism.
  • Ablation Tests: We confirm that the presence of literal matches significantly simplifies the task, enabling models to achieve high accuracy in answering questions. In contrast, when literal matches serve as distractors, they severely impair accuracy.
  • Chain-of-Thought (CoT) Prompting and Reasoning-based Models: While CoT prompting or reasoning-based models such as GPT-o1 (OpenAI et al., 2024) improve performance by encouraging step-by-step reasoning, they fail to fully mitigate the challenge, particularly in contexts exceeding 16K tokens.
@@ 2 Related Work
With the increasing popularity of long-context language modeling, numerous benchmarks have been introduced to evaluate this capability. Needle-in-a-Haystack (NIAH) is the most well-known and widely used benchmark (Mohtashami & Jaggi, 2023; Kamradt, 2023). However, due to performance saturation, various extensions have been proposed. These include increasing complexity by adding more needles, chaining needles to require inter-needle reasoning (fact-chaining), or incorporating arithmetic or code reasoning (Kamradt, 2023; Hsieh et al., 2024; Levy et al., 2024; Kuratov et al., 2024; Hengle et al., 2024; Zhang et al., 2024; Vodrahalli et al., 2024). Some tasks increase the complexity to such an extent that they become overly difficult even in short-context scenarios. For instance, BABILong includes tasks that perform poorly (e.g. the counting task achieves 28% accuracy) even without any irrelevant background text (0K) (Kuratov et al., 2024). Similarly, the Ancestral Tree Challenge (ATC) employs extensive fact-chaining, resulting in tasks that are overly complex even for short contexts ($<$1K) (Li et al., 2024). While such tasks challenge language models in long contexts, they raise the question of whether the tasks are inherently too complex for models to handle, regardless of context length.
Question Needles Keyword Types
Inv. $W_{n}$ is next to where [CHAR] lives. $W_{q}$ Countries, cities, states
Table 2: An example template of the proposed needle set in NoLiMa (all templates are available in Appendix A.) The placeholders [CHAR], $W_{q}$, and $W_{n}$ represent the randomly selected character (also the answer), the query keyword, and the needle keyword, respectively. Def.: default order. Inv.: inverted order.
@@
Another frequent pattern in many long-context benchmarks is the presence of literal matches between the facts required to answer a question and the question itself. This fact is not limited to synthetic recall-based tasks (e.g., vanilla NIAH, RULER retrieval-based sets) but also affects downstream-like QA-based benchmarks (Hsieh et al., 2024; Liu et al., 2024; Zhang et al., 2024; Bai et al., 2024; Yen et al., 2024), which often implicitly include literal matches between the relevant document and the question. Although many of these studies introduce complexity by adding similar documents as distractors, literal matches can still provide cues. These cues may help models focus on potential relevant facts based on matches, as attention mechanisms excel at recalling repetitive patterns (Olsson et al., 2022; Arora et al., 2024). We later demonstrate to what extent literal matches simplify recall-based questions (cf. 4.4.4). To quantify the prevalence of these matches in popular benchmarks, we compute ROUGE (R-1, R-2, and R-L) precision scores [^1] We use precision as our metric to measure how many of the question’s tokens exist in the relevant context, rather than the reverse. (Lin, 2004) between the needle (in recall-based tasks), the relevant document (in multi-document setups), or the full document (in long-document QA) and the corresponding question. This analysis measures the degree of literal overlap between the question and the context. Based on the scores in Table 1, NoLiMa demonstrates significantly less literal overlap compared to other datasets.
@@ 3 NoLiMa
To make NoLiMa an effective benchmark for evaluating LLM long-context abilities, we impose several constraints on the needle set. (i) We select keywords that ensure simplicity – so that, without irrelevant context, the associations are clear and the model can identify the correct answer. (ii) We randomize the assignment of character names from a diverse pool to minimize sensitivity to tokenization problems and mitigate ethnic bias (Navigli et al., 2023; Jiang et al., 2024). Names already occurring in the haystacks are excluded. (iii) We ensure $W_{n}$ is uniquely associated with $W_{q}$, avoid language-based cues and employ preface phrases to isolate needles from preceding context. See Appendix A for details.
Figure 1: Haystack conflicting information filtering pipeline
@@ 3.1 Haystack Filtering Pipeline
To ensure that the haystack does not contain: (1) Any distracting words that have extreme literal or high semantic similarities with the key points mentioned in the question (2) Any information that explicitly or in an inferrable case be a potential false answer to the question, we devise a filtering process.
@@
Conflicting Information Filtering.
In this step, we implement a semi-automatic redaction process to detect and remove such conflicting information. As shown in Figure 1, this process takes the haystack text—already filtered for distractors—along with questions from our needle set as input. Assuming the model should infer cases within short contexts, we scan the input texts in smaller chunks. [^2] With an 800-character stride and a 1000-character chunk size ($\sim$250 tokens). To identify potential answers within a chunk, we pair each question with the chunk and input them into an instruction-tuned language model, along with a short instruction and few-shot examples. The model responds with either “N/A” (indicating no relevant information was found) or an explanation identifying a possible conflict. Flagged examples are manually reviewed [^3] All manual reviews—in both filtering steps—were conducted by one of the authors. to determine whether the identified information should be removed. If no conflicts are found, the text remains unchanged. This process is repeated across all selected haystacks until no further removals are necessary.
@@ 4 Experiments
4.1 Dataset Configuration
In NoLiMa, we use 5 groups of needles, each with two fact-order variations: default and inverted. In the default order, the answer character always precedes the needle keyword $W_{n}$. In the inverted order, the fact is conveyed with the character name placed after $W_{n}$. Each group includes 2–6 keyword sets, with some sets containing multiple $W_{q}$ items to produce both one-hop and two-hop examples. This setup results in 58 question-needle pairs in total. To generate the haystacks, we select 10 open-licensed books, ensuring each covers at least 50K tokens. Using the filtering mechanism described in Section 3.1, we process the text to prepare it for haystack construction. To mitigate potential memorization issues—since these books are publicly available—we construct haystacks by concatenating short snippets. Specifically, we iteratively and randomly select a book, extract a continuous snippet (under 250 tokens), and append it to the haystack until it exceeds 2K lines, resulting in haystacks exceeding 60K tokens. In all experiments, each needle is placed 26 times at equal intervals across the evaluated context length. With 5 randomly generated haystacks, 58 question-needle pairs, and 26 placements per context length, this setup results in 7,540 tests per context length experiment.
@@ 4.1 Dataset Configuration
Command R+ 128K $<$1K 90.9 (77.3) 77.0 73.5 66.3 39.5 21.3 7.4
Mistral Large 2 128K 2K 87.9 (74.7) 86.1 85.5 73.3 51.5 32.6 18.7
Claude 3.5 Sonnet 200K 4K 87.6 (74.4) 85.4 84.0 77.6 61.7 45.7 29.8
@@
Table 3: NoLiMa benchmark results on the selected models. Following Hsieh et al. (2024), we report the effective length alongside the claimed supported context length for each model. However, we define the effective length as the maximum length at which the score remains above a threshold set at 85% of the model’s base score (shown in parentheses). Scores exceeding this threshold are underlined. Scores that are below 50% of the base score are shaded in red.
@@ 4.2 Models
For the filtering process, we opted using Llama 3.3 70b instruction tuned model (Meta, 2024). As a control test, for each question, we place its needle in 100 randomly selected chunks to verify whether the model (1) understands the filtering task and (2) is familiar with the facts and capable of inferring the answer. The model achieves a score of 99.8% in this test, indicating its ability to effectively flag conflicting information from the haystacks.
For the evaluation process, we select five closed-source models: GPT-4o, GPT-4o Mini (Hurst et al., 2024), Gemini 1.5 Pro, Flash (Team et al., 2024a), and Claude 3.5 Sonnet (Anthropic, 2024), along with seven open-weight Llama models: The Llama 3.x model family (3.1 8B, 70B, 405B, and 3.3 70B) (Dubey et al., 2024; Meta, 2024), Mistral Large (Mistral, 2024), Command R+ (Cohere For AI, 2024), and Jamba 1.5 Mini (Team et al., 2024b). All these models are well-known and widely used in long-context setups. In our analysis on reasoning-based prompting and models, we evaluate GPT-o1, GPT-o3 Mini (OpenAI et al., 2024; OpenAI, 2025), and DeepSeek-R1 Distill-Llama-70B (DeepSeek-AI et al., 2025). More details regarding model versions and deployment details are described in Appendix B.
@@ 4.3 Evaluation Setup & Metric
Models are evaluated on all tasks over context lengths of 250, 500, 1K, 2K, 4K, 8K, 16K, and 32K. To take into account how models would perform on NoLiMa regardless of long-context scenario, we control the difficulty of the task by by reporting a base score. Evaluations at context lengths of 250, 500, and 1K are used to compute the base score. These three are the shortest contexts. If a model can solve the task at these lengths, then any deterioration of its performance at greater lengths is expected to be solely due to its difficulties with generalizing over long contexts. For each question-needle example, we compute the average score over 5 haystacks, then take the maximum score of that example across the 250, 500, and 1K tests. The final base score is obtained by averaging these maximum scores across all question-needle examples. Inspired by Hsieh et al. (2024), we also report the effective length of each model. While they use the performance of Llama 2 model at a 4K length as a threshold (85.6%), we define the threshold as 85% of the base score. Thus, the effective length of a model is the largest tested length that exceeds this threshold. Additionally, some plots show the normalized score, calculated by dividing the accuracy score by the base score.

### v3

@@ Abstract
Recent large language models (LLMs) support long contexts ranging from 128K to 1M tokens. A popular method for evaluating these capabilities is the needle-in-a-haystack (NIAH) test, which involves retrieving a “needle” (relevant information) from a “haystack” (long irrelevant context). Extensions of this approach include increasing distractors, fact chaining, and in-context reasoning. However, in these benchmarks, models can exploit existing literal matches between the needle and haystack to simplify the task. To address this, we introduce NoLiMa, a benchmark extending NIAH with a carefully designed needle set, where questions and needles have minimal lexical overlap, requiring models to infer latent associations to locate the needle within the haystack. We evaluate 13 popular LLMs that claim to support contexts of at least 128K tokens. While they perform well in short contexts ($<$1K), performance degrades significantly as context length increases. At 32K, for instance, 11 models drop below 50% of their strong short-length baselines. Even GPT-4o, one of the top-performing exceptions, experiences a reduction from an almost-perfect baseline of 99.3% to 69.7%. Our analysis suggests these declines stem from the increased difficulty the attention mechanism faces in longer contexts when literal matches are absent, making it harder to retrieve relevant information. Even models enhanced with reasoning capabilities or CoT prompting struggle to maintain performance in long contexts. We publicly release the dataset and evaluation code at https://github.com/adobe-research/NoLiMa. [^1] Recent models not covered in the main body of this paper (such as GPT-4.1) have improved long-context performance. However, the main finding of this paper also holds for these newer models: performance on NoLiMa starts declining rapidly on contexts that are relatively short compared to their claimed context length. See Appendix E for detailed results.
@@ 1 Introduction
We evaluate NoLiMa over 13 state-of-the-art language models, all claiming to support token lengths of at least 128K, including GPT-4o, Gemini 1.5 Pro, and Llama 3.3 70B (Hurst et al., 2024; Gemini Team et al., 2024; Meta, 2024). Unlike NIAH-based evaluations, which contain literal matches and exhibit near-saturated performance, NoLiMa presents a more demanding challenge that highlights the limitations of these models. Their performance declines noticeably as context length increases, with considerable drops even at 2K–8K tokens. For instance, at 32K tokens, 11 out of 13 models achieve only half of their short-context performance.
@@
  • Impact of Latent Hops and Fact Direction: We demonstrate how the number of associative reasoning steps (latent hops) and the ordering of elements within a fact statement influence task performance. (§ 4.4.1)
  • Context Length vs. Needle Position: Our aligned-depth analysis shows that as latent reasoning complexity grows, performance depends more on context length than needle position. Without surface cues, longer contexts overwhelm the attention mechanism. (§ 4.4.2)
  • Chain-of-Thought (CoT) Prompting and Reasoning-based Models: While CoT prompting or reasoning-based models such as GPT-o1 (OpenAI et al., 2024) improve performance by encouraging step-by-step reasoning, they fail to fully mitigate the challenge, particularly in contexts exceeding 16K tokens. (§ 4.4.3)
  • Ablation Tests: We confirm that the presence of literal matches significantly simplifies the task, enabling models to achieve high accuracy in answering questions. In contrast, when literal matches serve as distractors, they severely impair accuracy. (§ 4.4.4)
@@ 2 Related Work
With the increasing popularity of long-context language modeling, numerous benchmarks have been introduced to evaluate this capability. Needle-in-a-Haystack (NIAH) is the most well-known and widely used benchmark (Mohtashami & Jaggi, 2023; Kamradt, 2023). However, due to performance saturation, various extensions have been proposed. These include increasing complexity by adding more needles, chaining needles to require inter-needle reasoning (fact-chaining), or incorporating arithmetic or code reasoning (Kamradt, 2023; Hsieh et al., 2024; Levy et al., 2024; Kuratov et al., 2024; Hengle et al., 2024; Zhang et al., 2024; Vodrahalli et al., 2024). Some tasks increase the complexity to such an extent that they become overly difficult even in short-context scenarios. For instance, BABILong includes tasks that perform poorly (e.g., the counting task achieves 28% accuracy) even without any irrelevant background text (0K) (Kuratov et al., 2024). Similarly, the Ancestral Tree Challenge (ATC) employs extensive fact-chaining, resulting in tasks that are overly complex even for short contexts ($<$1K) (Li et al., 2024). While such tasks challenge language models in long contexts, they raise the question of whether the tasks are inherently too complex for models to handle, regardless of context length.
Question Needle Keyword Type
Which character has been to $W_{q}$? (Def.) Actually, [CHAR] lives next to the $W_{n}$. $W_{n}$ Buildings&Landmarks (e.g., Semper Opera)
(Inv.) $W_{n}$ is next to where [CHAR] lives. $W_{q}$ Countries, cities, states (e.g., Dresden)
Table 2: An example template of the proposed needle set in NoLiMa. (All templates are available in Appendix A.) The placeholders [CHAR], $W_{q}$, and $W_{n}$ represent the randomly selected character (also the answer), the query keyword, and the needle keyword, respectively. Def.: default order. Inv.: inverted order.
@@
Another frequent pattern in many long-context benchmarks is the presence of literal matches between the facts required to answer a question and the question itself. This fact is not limited to synthetic recall-based tasks (e.g., vanilla NIAH, RULER retrieval-based sets) but also affects downstream-like QA-based benchmarks (Hsieh et al., 2024; Liu et al., 2024; Zhang et al., 2024; Bai et al., 2024; Yen et al., 2024), which often implicitly include literal matches between the relevant document and the question. Although many of these studies introduce complexity by adding similar documents as distractors, literal matches can still provide cues. These cues may help models focus on potential relevant facts based on matches, as attention mechanisms excel at recalling repetitive patterns (Olsson et al., 2022; Arora et al., 2024). We later demonstrate to what extent literal matches simplify recall-based questions (cf. 4.4.4). To quantify the prevalence of these matches in popular benchmarks, we compute ROUGE (R-1, R-2, and R-L) precision scores [^2] We use precision as our metric to measure how many of the question’s tokens occur in the relevant context, rather than the reverse. (Lin, 2004) between the question and the context – the needle (in recall-based tasks), the relevant document (in multi-document setups), or the full document (in long-document QA). This analysis measures the degree of literal overlap between the question and the context. Table 1 demonstrates that NoLiMa has much less literal overlap than other datasets.
@@ 3 NoLiMa
To make NoLiMa an effective benchmark for evaluating LLM long-context abilities, we impose several constraints on the needle set. (i) We select keywords that ensure simplicity – so that, without irrelevant context, the associations are clear and the model can identify the correct answer. (ii) We randomize the assignment of character names from a diverse pool to minimize sensitivity to tokenization problems and mitigate ethnic bias (Navigli et al., 2023; Jiang et al., 2024). Names already occurring in the haystacks are excluded. (iii) We ensure $W_{n}$ is uniquely associated with $W_{q}$, avoid language-based cues, and in most cases employ preface phrases–short lead-ins or contextual buildup (e.g., “Actually,” “In 2013, after waiting in line…”)–to isolate needles from preceding context. See Appendix A for details.
Figure 1: Haystack filtering pipeline for undesired or misleading content
@@ 3.1 Haystack Filtering Pipeline
We devise a filtering process to ensure that the haystack does not contain: (1) Any distracting words that have extreme literal or high semantic similarity with the key points mentioned in the question, (2) Any information that explicitly or through inference is a potential false answer to the question.
@@
Filtering Undesired Answer Candidates.
In this step, we implement a semi-automatic redaction process to detect and remove text spans that could be interpreted as plausible but unintended answers. As shown in Figure 1, this process takes the haystack text—already filtered for distractors—along with questions from our needle set as input. Assuming the model should infer cases within short contexts, we scan the input texts in smaller chunks. [^3] With an 800-character stride and a 1000-character chunk size ($\sim$250 tokens). To identify potential answers within a chunk, we pair each question with the chunk and input them into an instruction-tuned language model, along with a short instruction and few-shot examples. The model responds with either “N/A” (indicating no relevant information was found) or an explanation identifying a possible candidate answer. Flagged examples are manually reviewed [^4] All manual reviews—in both filtering steps—were conducted by one of the authors. to determine whether the identified content should be removed. If no relevant content is identified, the text remains unchanged. This process is repeated across all selected haystacks until no further removals are necessary.
@@ 4 Experiments
Command R+ 128K $<$1K 90.9 (77.3) 77.0 73.5 66.2 39.5 21.3 7.4
Gemini 2.0 Flash 1M 4K 89.4 (76.0) 87.7 87.5 77.9 64.7 48.2 41.0
Mistral Large 2 128K 2K 87.9 (74.7) 86.1 85.5 73.3 51.4 32.6 18.8
Claude 3.5 Sonnet 200K 4K 87.5 (74.4) 85.4 84.0 77.6 61.7 45.7 29.8
@@
Table 3: NoLiMa benchmark results on the selected models. Following Hsieh et al. (2024), we report the effective length alongside the claimed supported context length for each model. However, we define the effective length as the maximum length at which the score remains above a threshold, set at 85% of the model’s base score (shown in parentheses). Scores exceeding this threshold are underlined. Scores that are below 50% of the base score are shaded in red.
4.1 Dataset Configuration
In NoLiMa, we use 5 groups of needles, each with two “word order” variations: default and inverted. In the default order, the answer character (CHAR) precedes the needle keyword ($W_{n}$), following the pattern “…[CHAR] …$W_{n}$” (see Table 2, column “Needle”). In the inverted order, the character name follows $W_{n}$, yielding the pattern “$W_{n}$[CHAR] …”. Each group includes 2–6 keyword sets, with some sets containing multiple $W_{q}$ items to produce both one-hop and two-hop examples. This setup results in 58 question-needle pairs in total. To generate the haystacks, we select 10 open-license books, ensuring each covers at least 50K tokens. Using the filtering mechanism described in Section 3.1, we process the text to prepare it for haystack construction. To mitigate potential memorization issues—since these books are publicly available—we construct haystacks by concatenating short snippets. Specifically, we iteratively and randomly select a book, extract a continuous snippet (under 250 tokens), and append it to the haystack until it exceeds 2K lines, resulting in haystacks exceeding 60K tokens. In all experiments, each needle is placed 26 times at equal intervals across the evaluated context length. With 5 randomly generated haystacks, 58 question-needle pairs, and 26 placements per context length, this setup results in 7,540 tests per context length experiment.
@@ 4.2 Models
For the filtering process, we opted for using the Llama 3.3 70b instruction tuned model (Meta, 2024). As a control test, for each question, we place its needle in 100 randomly selected chunks to determine whether the model (1) understands the filtering task and (2) is familiar with the facts and capable of inferring the answer. The model achieves a score of 99.8% in this test, indicating its ability to effectively flag conflicting information from the haystacks.
For the evaluation process, we select five closed-source models: GPT-4o, GPT-4o Mini (Hurst et al., 2024), Gemini 1.5 Pro Flash, Gemini 2.0 Flash (Gemini Team et al., 2023, 2024) and Claude 3.5 Sonnet (Anthropic, 2024), along with seven open-weight Llama models: The Llama 3.x model family (3.1 8B, 70B, 405B, and 3.3 70B) (Dubey et al., 2024; Meta, 2024), Mistral Large (Mistral, 2024), Command R+ (Cohere For AI, 2024), and Jamba 1.5 Mini (Team et al., 2024). All these models are well-known and widely used in long-context setups. In our analysis on reasoning-based prompting and models, we evaluate GPT-o1, GPT-o3 Mini (OpenAI et al., 2024; OpenAI, 2025), and DeepSeek-R1 Distill-Llama-70B (DeepSeek-AI et al., 2025). More details regarding model versions and deployment details are described in Appendix B.
@@ 4.3 Evaluation Setup & Metric
Models are evaluated on all tasks over context lengths of 250, 500, 1K, 2K, 4K, 8K, 16K, and 32K. To take into account how models would perform on NoLiMa regardless of long-context scenario, we control the difficulty of the task by reporting a base score. Evaluations at context lengths of 250, 500, and 1K are used to compute the base score. These three are the shortest contexts. If a model can solve the task at these lengths, then any deterioration of its performance at greater lengths is expected to be solely due to its difficulties with generalizing over long contexts. For each question-needle example, we compute the average scores over 5 haystacks, then take its maximum score across the 250, 500, and 1K tests. The final base score is obtained by averaging these maximum scores across all question-needle examples. Inspired by Hsieh et al. (2024), we also report the effective length of each model. While they use the performance of Llama 2 at 4K context length as a threshold (85.6%), we define the threshold as 85% of the base score. Thus, the effective length of a model is the largest tested length that exceeds this threshold. Additionally, some plots show the normalized score, calculated by dividing the accuracy score by the base score.
4.4 Results
Table 3 presents the performance results of all NoLiMa tests on the selected models. Most models achieve high base scores, indicating that the designed needle set is relatively simple to answer in shorter contexts. Even models with base scores exceeding 90.0% exhibit a significantly shorter effective length than their claimed lengths, generally limited to $\leq$2K tokens, with GPT-4o being an exception. While GPT-4o demonstrates strong overall performance, it fails to generalize effectively beyond 8K tokens.1 Out of the 13 models, 11 exhibit performance at 32K lengths that is half or less of their base scores. For comparison, in other benchmarks with similar settings, such as BABILong (QA1) (Kuratov et al., 2024) and RULER (Hsieh et al., 2024), Llama 3.1 70B achieves effective lengths of 16K [^5] In BABILong, the effective length is also based on 85% of the 0K base performance threshold and 32K, respectively. However, in NoLiMa, Llama 3.1 70B has an effective length of only 2K and shows a significant drop in performance at 32K lengths (42.7% vs. 94.3% base score). Models such as Claude 3.5 Sonnet, Gemini 1.5 Flash, GPT-4o mini, and Llama 3.1 8B may have weaker base scores, but their effective lengths are calculated relative to these scores. This reveals an interesting observation: Models like Claude 3.5 Sonnet and Gemini 2.0 Flash, despite having a lower base score, may underperform in shorter contexts but demonstrate better length generalization than models with higher base scores, such as Llama 3.1 70B and Llama 3.3 70B. In fact, both Sonnet and Gemini 2.0 Flash achieve even higher raw scores in 4K-token experiments compared to some higher-base-score models.
Based on the dataset construction method described in Section 4.1, NoLiMa can generate haystacks at any desired length. We applied this to test GPT-4o and Gemini 2.0 Flash at 64K and 128K tokens. GPT-4o maintained over 50% of its base score at 128K, while Gemini 2.0 Flash dropped to just 16.4%. Full results and evaluation details are provided in Appendix E.
(a) One-hop vs. Two-hop
(b) Default vs. Inverted
Figure 2: Impact of (a) number of hops and (b) inversion of order (“[CHAR] …$W_{n}$” vs. “$W_{n}$CHAR”) on normalized performance across GPT-4o and Llama 3.3 70B models. The red dotted line indicates the 0.85 effective threshold.
4.4.1 Latent hops & Inversion
As discussed in Section 3, our needle set also includes examples requiring two-hop associative linking from the question keyword to the needle keyword. To evaluate the impact on length generalization, Figure 2(a) presents the normalized performance of two top-performing models on one-hop and two-hop tasks. It is evident that, for the same context lengths, questions involving two-hop latent reasoning steps are more challenging than those requiring one-hop reasoning. Notably, the performance gap between one-hop and two-hop tasks widens with increasing context lengths. GPT-4o demonstrates impressive generalization performance, handling both types of examples effectively even at context lengths up to 4K. A detailed breakdown of performance on one-hop and two-hop examples is provided in Appendix F, complementing the aggregate results shown in Table 3.
Each group of needles includes both a default and an inverted template and Figure 2(b) shows that inverted examples are more challenging to answer. We argue this difficulty arises from the model’s causal attention mechanism, particularly in longer contexts where attention signals weaken. In the default template, the question – in particular $W_{q}$ – can link directly to $W_{n}$, which generally will contain information about the character’s name since the name appears earlier in the sequence. This allows the model to backtrace effectively from $W_{q}$ through $W_{n}$ to the character. In the inverted template, $W_{q}$ may still attend to $W_{n}$, but since the fact is incomplete (the character hasn’t been mentioned yet), the model cannot use that attention to resolve the question. Instead, it must rely on weaker signals encoded in the character’s name to establish the link, which becomes harder with longer contexts due to diminishing attention strength. While these findings shed light on the challenge, deeper mechanistic analysis is beyond the scope of this paper and requires further study.